OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

you just want the answer?

OpenStudy (anonymous):

y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2) yes please i need to finish up quick

OpenStudy (anonymous):

are you sure those are the answers?

OpenStudy (anonymous):

cause slope is 1/4

OpenStudy (anonymous):

its the answers give, im confused by the teaching on this site the answers seem bogus

OpenStudy (anonymous):



OpenStudy (anonymous):

thats the answer unless the points you gave are wrong

OpenStudy (anonymous):

my online class is teaching us the wrong thing lol thanks anyway. ive searched everywhere for this answer and no one seems to give any of those listed